OpenStudy (anonymous):

Ron is twice as old as his nephew. The sum of their ages is 108. How old is Ron

OpenStudy (shayaan_mustafa):

if nephew is x then ron is 2x (as ron is twice as old as his nephew), right? then equation will be. x+2x=108 3x=108 x=108/3 x=36

OpenStudy (anonymous):

so then ron is 36?

OpenStudy (anonymous):

No ron is 72 years old since ron is twice as old.

OpenStudy (shayaan_mustafa):

no no

OpenStudy (shayaan_mustafa):

x=36 is the age of nephew twice of nephew is ron's age i.e. 2x=2(36) 2x=72